# Image Segmentation in Foundation Model Era: A Survey

### v1

@@ Abstract
Image segmentation is a long-standing challenge in computer vision, studied continuously over several decades, as evidenced by seminal algorithms such as N-Cut, FCN, and MaskFormer. With the advent of foundation models (FMs), contemporary segmentation methodologies have embarked on a new epoch by either adapting FMs (e.g., CLIP, Stable Diffusion, DINO) for image segmentation or developing dedicated segmentation foundation models (e.g., SAM). These approaches not only deliver superior segmentation performance, but also herald newfound segmentation capabilities previously unseen in deep learning context. However, current research in image segmentation lacks a detailed analysis of distinct characteristics, challenges, and solutions associated with these advancements. This survey seeks to fill this gap by providing a thorough review of cutting-edge research centered around FM-driven image segmentation. We investigate two basic lines of research – generic image segmentation (i.e., semantic segmentation, instance segmentation, panoptic segmentation), and promptable image segmentation (i.e., interactive segmentation, referring segmentation, few-shot segmentation) – by delineating their respective task settings, background concepts, and key challenges. Furthermore, we provide insights into the emergence of segmentation knowledge from FMs like CLIP, Stable Diffusion, and DINO. An exhaustive overview of over 300 segmentation approaches is provided to encapsulate the breadth of current research efforts. Subsequently, we engage in a discussion of open issues and potential avenues for future research. We envisage that this fresh, comprehensive, and systematic survey catalyzes the evolution of advanced image segmentation systems.
@@ 1 Introduction
Related Surveys and Differences. In the past decade, many surveys have studied image segmentation from various perspectives. For example, [73] reviews region- and boundary-based segmentation methods in 2015. With the transition to the deep learning era, a series of works [74, 70, 75, 76, 77, 78] has summarized progress in generic segmentation tasks like semantic, instance and panoptic segmentation. A recent study [79] focuses on the specific task of open-vocabulary segmentation. Other studies delve into crucial aspects of image segmentation, such as evaluation protocols [80] or loss functions [81]. In addition, there exist surveys that focus on segmentation techniques in specialized domains, e.g., video [82], medical imaging [83, 84].
@@
Survey Organization. The reminder of this paper is structured as follows. §2 presents essential background on image segmentation and FMs. §3 highlights the emergency of segmentation knowledge from existing FMs. §4 and §5 review the most important FM-based image segmentation methods, mainly from the past three years. §6 raises open issues and future directions. We conclude the paper in §7.
§4§5§4.1§4.2§4.3§5.1§5.2§5.3[87][88][89][90][91][92][93][94][95][96][97][98][99][100][101][102][103][104][105][106][107][108][109][50][110][111][112][113][114][115][59][60][116][117][118][119][120][121][122][123][124]
Figure 1: The taxonomy based on image segmentation tasks and foundation models. Typical solutions are shown for each category.
@@ 2.1.2 Image Segmentation Category
According to whether $P$ is provided, we categorize image segmentation methods into two classes (Fig. 1): generic image segmentation (GIS) and promptable image segmentation (PIS).
@@ 2.2.1 Language Foundation Model
$\bullet$ Multimodal Large Language Models (MLLMs). MLLMs [72] are multimodal extensions of LLMs by bringing together the reasoning power of LLMs with the capability to process non-textual modalities (e.g., vision, audio). This kind of unification expands the impact of LLMs with novel interfaces and capabilities, enabling them to solve new and more complicated tasks beyond the NLP domain. MLLMs represent the next level of LLMs. On one hand, multimodal perception is a natural way for knowledge acquisition and interaction with the real world, and thus serves as a fundamental component for achieving AGI; on the other hand, the multimodal extension expands the potential of pure language modeling to more complex tasks in, e.g., robotics and autonomous driving.
@@ 2.2.2 Visual Foundation Model
$\small\mathcal{L}_{i2t}=-\log\left[\frac{\exp(\text{sim}(\bm{x}_{k},\bm{t}_{k})/\tau)}{\sum_{j=1}^{J}\exp(\text{sim}(\bm{x}_{k},\bm{t}_{j})/\tau)}\right].$ (2)
@@
$\small\bm{z}_{t}\sim\mathcal{N}\left(\sqrt{\alpha_{t}}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}\right),$ (3)
$\small\mathcal{L}_{\text{DM}}\!=\!\mathbb{E}_{\bm{z}_{0},\epsilon\sim\mathcal{N}(0,1),t}\left[||\epsilon-\epsilon_{\theta}(\bm{z}_{t}(\bm{z}_{0},\epsilon),t;\mathcal{C})||_{2}^{2}\right],$ (4)
$\small\mathcal{L}_{\text{LDM}}\!=\!\mathbb{E}_{\mathcal{E}(\bm{z}_{0}),\epsilon\sim\mathcal{N}(0,1),t}\left[||\epsilon-\epsilon_{\theta}(\bm{z}_{t}(\mathcal{E}(\bm{z}_{0}),\epsilon),t;\mathcal{C})||_{2}^{2}\right].$ (5)
$\bullet$ Segment Anything (SAM). SAM [49] has sparked a revolution in the field of image segmentation, and profoundly influences the development of large, general-purposed models in computer vision. Unlike the aforementioned vision FMs, SAM is built specifically for segmentation. Particularly, it is a promptable model trained on a corpus of 1 billion masks from 11 million images using a promptable segmentation task. It achieves powerful zero-shot task generality to tackle various image segmentation tasks, and allows the use of “prompts” in forms of points, masks, boxes, even language to enhance segmentation interactivity. Beyond this, SAM has shown promising capabilities in a multitude of tasks, including medical imaging [111], image editing [134], video segmentation [135]. We refer readers to [86] for a comprehensive understanding of SAM’s usage scenarios.
@@ 3.1 Segmentation Emerges from CLIP
$\small\text{SelfAttention}(\bm{q},\bm{k},\bm{v})=\underbrace{\text{softmax}\left(\bm{q}\bm{k}^{\top}/\sqrt{d}\right)}_{\text{self-attention matrix}}\bm{v},$ (6)
@@ 4 Foundation Model based GIS
This section presents a comprehensive review of recent advances in FM-based GIS, including semantic (§4.1), instance (§4.2) and panoptic segmentation (§4.3). Our discussions are approached from a technical perspective to elucidate the fundamental concepts and highlight the roles of FMs in GIS.
@@ 4.1.2 DM-based Solution
$\bullet$ T2I-DMs as Semantic Segmentation Data Synthesizer. Collecting and annotating images with pixel-wise labels is time-consuming and laborious, and thus always a challenge to semantic segmentation. With recent advancements in AIGC, many studies [141, 196, 99, 98] explore the potential of T2I-DMs to build large-scale segmentation dataset (including synthetic images and associated mask annotations), which can serve as a more cost-effective data source to train any existing semantic segmentation models. The idea has also been adopted in specialist domains like medical image segmentation [197]. Rather than directly generating synthetic masks, some works [198, 199, 200] employ T2I-DMs for data augmentation based on a few labeled images.
@@ 4.3.1 CLIP-based Solution
$\bullet$ CLIP as Zero-Shot Mask Classifier. Most recent panoptic segmentation approaches [104, 96, 220, 97, 212, 105, 221, 211] follow the query-based mask classification framework introduced in MaskFormer [22] / Mask2Former [23]. They generate class-agnostic mask proposals first and then utilize CLIP to classify the proposals, thereby empowering MaskFormer and Mask2Former open-vocabulary segmentation capabilities. MaskCLIP [104] introduces a set of mask class tokens to extract mask representations more efficiently. MasQCLIP [96] augments MaskCLIP by applying additional projections to mask class tokens to obtain optimal attention weights. OPSNet [97] learns more generalizable mask representations based on CLIP visual encoder that are subsequently used to enhance query embeddings. Unpair-Seg [105] presents a weakly supervised framework that allows the model to benefit from cheaper image-text pairs. It learns a feature adapter to align mask representations with text embeddings, which are extracted from CLIP’s visual and language encoders respectively. Despite the advances, these methods still require training a separate model for each task to achieve the best performance. Freeseg [221] and DaTaSeg [222] design all-in-one models with the same architecture and inference parameters to establish remarkable performance in open-vocabulary semantic, instance, and panoptic segmentation problems. OMG-Seg [223] introduces a unified query representation to unify different task outputs, and is able to handle ten segmentation tasks across different datasets.
@@ 4.3.4 SAM-based Solution
$\bullet$ Towards Semantic-Awareness SAM. While SAM demonstrates strong zero-shot performance, it produces segmentation without semantic meaning. This drives many research efforts, e.g., Semantic-SAM [109], SEEM [50], to enhance the semantic-awareness of SAM. In addition to visual prompts in SAM for interactive segmentation, these models learn generic object queries to achieve generic segmentation in both semantic and instance levels. In addition, the models are generally trained on a combination of multiple datasets with semantic annotations, such as COCO [229], ADE20K [230], PASCAL VOC [231].
@@ 5.1.1 SAM-based Solution
$\bullet$ Multi-Granularity Interactive Segmentation. Most existing interactive segmentation methods determines a single segmentation mask based on users’ input, which ignores spatial ambiguity. In contrast, SAM introduces a multi-granularity interactive segmentation pipeline, i.e., for each user interaction, desired segmentation region may be the concept of objects with different parts nearby. To improve the segmentation quality, HQ-SAM [218] proposes a lightweight high-quality output token replace the original SAM’s output token. After training on 44K highly-accurate masks, HQ-SAM significantly boosts the mask prediction quality of SAM. Since SAM is class-agnostic, a line of works [232, 233] tunes SAM by aligning the query-segmented regions with corresponding textual representations from CLIP, while [109] designs a SAM-like framework that supports multi-granularity segmentation using the captioned SAM data. Although these multi-granularity interactive segmentation approaches alleviate spatial ambiguity, they result in excessive output redundancy and limited scalability. To solve this, GraCo [110] explores granularity-controllable interactive segmentation, which allows precise control of prediction granularity to resolve ambiguity.
@@ 5.2.2 DM-based Solution
Figure 2: MLLMs driven solutions lead to more powerful pixel reasoning and understanding capabilities, e.g., multi-target reasoning segmentation, instance segmentation with text descriptions, referring segmentation and conversation. (Figure adapted courtesy of [60])
@@ 5.2.3 LLMs/MLLMs-based Solution
The success of LLMs/MLLMs has showcased incredible reasoning ability and can answer complex questions, thereby bringing new possibilities to achieve new pixel reasoning and understanding capabilities (Fig. 2). In particularly, LISA [59] studies a new segmentation task, called reasoning segmentation. Different from traditional referring segmentation, the segmentors in this setting are developed to segment the object based on implicit query text involving complex reasoning. Notably, the query text is not limited to a straightforward reference (e.g., “the front-runner”), but a a more complicated description involving complex reasoning or world knowledge (e.g., “who will win the race?”). LISA employs LLaVA [258] to output a text response based on the input image, text query, and a [seg] token. The embedding for the customized [seg] token is decoded into the segmentation mask via SAM decoder. Afterwards, LISA++ [259] promotes LISA to differentiate individuals within the same category and enables more natural conversation in multi-turn dialogue. Based on these works, many efforts have been devoted to promote the reasoning capability and segmentation accuracy. LLM-Seg [260] proposes using SAM to generate a group of mask proposals that selects the best-suited answer as the final segmentation prediction. Next-Chat [261] adds a [trigger] token that depicts the coordinate of the object box as a supplementary input for MLLM to help generate better masks. Similarly, GSVA [262] introduces a rejection token [rej] to relieve the empty-target case where the object referred to in the instructions does not exist in the image, leading to the false-positive prediction. Except for the functional token incorporation, [263, 264] propose using diverse textual descriptions, such as object attribute and part, to enhance the object-text connection for accurate reasoning results. Regarding reasoning costing, PixelLLM [60] introduces a lightweight decoder to reduce the computational cost in the reasoning process. Osprey [265] extends MLLMs by incorporating fine-grained mask regions into language instruction, and delivers remarkable pixel-wise visual understanding capabilities.
@@ 5.3.5 LLM/MLLM-based Solution
Figure 3: Illustration of VPImpainting [302], which solves image segmentation as a visual “fill-in-the-blank” problem. (Figure courtesy of [302])
@@ 5.3.6 In-Context Segmentation
The rapid progress of LLMs leads to an emergent capacity to learn in-context from given a handful of examples [38, 45]. Inspired by such an amazing pattern, some researchers aim to excavate a new and similar setting in computer vision, i.e., in-context segmentation (ICS). With the aim to segment a query image based on supports, ICS can be regarded as a sub-task of FSS. Yet, ICS requires no parameter updating and can be directly performed on pre-trained models without task-specific finetuning. The ICL-emerged LLMs mostly are generative models trained with masked language modelling or next token prediction strategy. Therefore, most works, through intuitive imitation, address ICS by exploring similar self-supervised patterns on visual models. VPImpainting [302] is a pioneering work that solves visual in-context learning as image inpainting. The architecture is illustrated in Fig. 3. It defines visual prompt as a grid-like single image containing an input-output example(s) and a query, then trains an inpainting model (via MAE [303]) to predict the rest of the image such that it is consistent with given example(s). With this basis, [304, 305, 306] propose to retrieve the most suitable examples from a large dataset as the support. Furthermore, Painter [307] and SegGPT [51] are vision generalists built on in-context learning. They unify various vision tasks into the in-context learning framework by carefully redefining outputs of core vision tasks into the same format of images. Some other works [308, 309] focus on establishing large vision model by formatting images, like language tokenizer, to a group of sequence as visual sentences, and then perform LLM-like training via next token prediction. It is worth noting that such a cultivation of visual autoregressive models require hundreds of billions of vision samples from varied vision tasks, e.g., image segmentation, depth estimation. PromptDiffusion [310] explores in-context learning for diffusion models by fine-tuning SD to generate the query mask conditioned on the support image-mask pair and the query image. Matcher [311] utilizes DINOv2 to locate the target in query images by bidirectional matching, and leverages the coarse location information as the prompts of SAM for segmentation. Tyche [312] extends ICS into a probabilistic segmentation framework by explicitly modeling training and testing uncertainty, and shows promising performance in medical image segmentation.

### v2

@@ Abstract
Image segmentation is a long-standing challenge in computer vision, studied continuously over several decades, as evidenced by seminal algorithms such as N-Cut, FCN, and MaskFormer. With the advent of foundation models (FMs), contemporary segmentation methodologies have embarked on a new epoch by either adapting FMs (e.g., CLIP, Stable Diffusion, DINO) for image segmentation or developing dedicated segmentation foundation models (e.g., SAM, SAM2). These approaches not only deliver superior segmentation performance, but also herald newfound segmentation capabilities previously unseen in deep learning context. However, current research in image segmentation lacks a detailed analysis of distinct characteristics, challenges, and solutions associated with these advancements. This survey seeks to fill this gap by providing a thorough review of cutting-edge research centered around FM-driven image segmentation. We investigate two basic lines of research – generic image segmentation (i.e., semantic segmentation, instance segmentation, panoptic segmentation), and promptable image segmentation (i.e., interactive segmentation, referring segmentation, few-shot segmentation) – by delineating their respective task settings, background concepts, and key challenges. Furthermore, we provide insights into the emergence of segmentation knowledge from FMs like CLIP, Stable Diffusion, and DINO. An exhaustive overview of over 300 segmentation approaches is provided to encapsulate the breadth of current research efforts. Subsequently, we engage in a discussion of open issues and potential avenues for future research. We envisage that this fresh, comprehensive, and systematic survey catalyzes the evolution of advanced image segmentation systems.
@@ 1 Introduction
§1§2§2.1§2.2§3§3.1§3.2§3.3§4§4.1§4.2§4.3§5§5.1§5.2§5.3§6§7
Figure 1: Overview of this survey.
Related Surveys and Differences. In the past decade, many surveys have studied image segmentation from various perspectives. For example, [73] reviews region- and boundary-based segmentation methods in 2015. With the transition to the deep learning era, a series of works [74, 70, 75, 76, 77, 78] has summarized progress in generic segmentation tasks like semantic, instance and panoptic segmentation. A recent study [79] focuses on the specific task of open-vocabulary segmentation, while [80] only studies Transformer-based segmentation. Other studies delve into crucial aspects of image segmentation, such as evaluation protocols [81] or loss functions [82]. In addition, there exist surveys that focus on segmentation techniques in specialized domains, e.g., video [83], medical imaging [84, 85].
@@
Survey Organization. Fig. 1 shows the structure of this survey. Section §2 presents essential background on image segmentation and FMs. §3 highlights the emergency of segmentation knowledge from existing FMs. §4 and §5 review the most important FM-based image segmentation methods, mainly from the past three years. §6 raises open issues and future directions. We conclude the paper in §7.
@@ 2.1.2 Image Segmentation Category
According to whether $P$ is provided, we categorize image segmentation methods into two classes: generic image segmentation (GIS) and promptable image segmentation (PIS).
@@ 2.2.1 Language Foundation Model
$\bullet$ Multimodal Large Language Models (MLLMs). MLLMs [72] are multimodal extensions of LLMs by bringing together the reasoning power of LLMs with the capability to process non-textual modalities (e.g., vision, audio). MLLMs represent the next level of LLMs. On one hand, multimodal perception is a natural way for knowledge acquisition and interaction with the real world, and thus serves as a fundamental component for achieving AGI; on the other hand, the multimodal extension expands the potential of pure language modeling to more complex tasks in, e.g., robotics and autonomous driving.
@@ 2.2.2 Visual Foundation Model
$\small\mathcal{L}_{i2t}=-\log\left[\frac{\exp(\text{sim}(\bm{x}_{k},\bm{t}_{k})/\tau)}{\sum_{j=1}^{J}\exp(\text{sim}(\bm{x}_{k},\bm{t}_{j})/\tau)}\right].\vspace{-3pt}$ (2)
@@
$\small\bm{z}_{t}\sim\mathcal{N}\left(\sqrt{\alpha_{t}}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}\right),\vspace{-3pt}$ (3)
$\small\mathcal{L}_{\text{DM}}\!=\!\mathbb{E}_{\bm{z}_{0},\epsilon\sim\mathcal{N}(0,1),t}\left[||\epsilon-\epsilon_{\theta}(\bm{z}_{t}(\bm{z}_{0},\epsilon),t;\mathcal{C})||_{2}^{2}\right],\vspace{-3pt}$ (4)
$\small\mathcal{L}_{\text{LDM}}\!=\!\mathbb{E}_{\mathcal{E}(\bm{z}_{0}),\epsilon\sim\mathcal{N}(0,1),t}\left[||\epsilon-\epsilon_{\theta}(\bm{z}_{t}(\mathcal{E}(\bm{z}_{0}),\epsilon),t;\mathcal{C})||_{2}^{2}\right].\vspace{-3pt}$ (5)
@@
$\bullet$ Segment Anything (SAM). SAM [49] has sparked a revolution in the field of image segmentation, and profoundly influences the development of large, general-purposed models in computer vision. Unlike the aforementioned vision FMs, SAM is built specifically for image segmentation, which is trained on a corpus of 1 billion masks from 11 million images using a promptable segmentation task. It achieves powerful zero-shot task generality to handle a wide range of image segmentation tasks, and allows for enhanced interactivity in segmentation through the use of “prompts” in forms of points, masks, boxes, and even language. Beyond this, SAM has shown promising capabilities in a multitude of tasks, including medical imaging [97], image editing [98], video segmentation [99]. Despite its capabilities, one downside of SAM is the computational expense associated with its heavy image encoder. However, SAM2 [100] addresses this by instead using an MAE pre-trained Hiera as the image encoder, yielding real-time speed and improved segmentation accuracy.
@@ 3.1 Segmentation Emerges from CLIP
$\small\text{SelfAttention}(\bm{q},\bm{k},\bm{v})=\underbrace{\text{softmax}\left(\bm{q}\bm{k}^{\top}/\sqrt{d}\right)}_{\text{self-attention matrix}}\bm{v},\vspace{-3pt}$ (6)
@@ 4 Foundation Model based GIS
This section presents a comprehensive review of recent advances in FM-based GIS, including semantic (§4.1), instance (§4.2) and panoptic segmentation (§4.3), as illustrated in Fig. 2. Our discussions are approached from a technical perspective to elucidate the fundamental concepts and highlight the roles of FMs in GIS.
@@ 4.1.2 DM-based Solution
$\bullet$ T2I-DMs as Semantic Segmentation Data Synthesizer. Collecting and annotating images with pixel-wise labels is time-consuming and laborious, and thus always a challenge to semantic segmentation. With recent advances in AIGC, many studies [106, 167, 168, 169] explore the potential of T2I-DMs to build large-scale segmentation dataset (including synthetic images and associated mask annotations), which serve as a more cost-effective data source to train any existing semantic segmentation models. The idea has also been adopted in specialist domains like medical image segmentation [170]. Rather than directly generating synthetic masks, some works [171, 172, 173] employ T2I-DMs for data augmentation based on a few labeled images.
@@ 4.3.1 CLIP-based Solution
$\bullet$ CLIP as Zero-Shot Mask Classifier. Most recent panoptic segmentation approaches [188, 189, 204, 191, 192, 205, 206, 190] follow the query-based mask classification framework introduced in MaskFormer [22] / Mask2Former [23]. They generate class-agnostic mask proposals first and then utilize CLIP to classify the proposals, thereby empowering MaskFormer and Mask2Former open-vocabulary segmentation capabilities. MaskCLIP [188] introduces a set of mask class tokens to extract mask representations more efficiently. MasQCLIP [189] augments MaskCLIP by applying additional projections to mask class tokens to obtain optimal attention weights. OPSNet [191] learns more generalizable mask representations based on CLIP visual encoder that are subsequently used to enhance query embeddings. Unpair-Seg [205] presents a weakly supervised framework that allows the model to benefit from cheaper image-text pairs. It learns a feature adapter to align mask representations with text embeddings, which are extracted from CLIP’s visual and language encoders respectively. Despite the advances, these methods still require training a separate model for each task to achieve the best performance. Freeseg [206] and DaTaSeg [207] design all-in-one models with the same architecture and inference parameters to establish remarkable performance in open-vocabulary semantic, instance, and panoptic segmentation problems. OMG-Seg [208] introduces a unified query representation to unify different task outputs, and is able to handle 10 segmentation tasks across different datasets.
@@ 4.3.4 SAM-based Solution
$\bullet$ Towards Semantic-Aware SAM. While SAM shows strong zero-shot performance, its outputs are semantic-agnostic. This drives many research efforts, e.g., Semantic-SAM [217], SEEM [50], to enhance the semantic-awareness of SAM. In addition to visual prompts in SAM for interactive segmentation, these models learn generic object queries to achieve generic segmentation in both semantic and instance levels. In addition, the models are generally trained on a combination of multiple datasets with semantic annotations, such as COCO [218], ADE20K [219], PASCAL VOC [220].
Figure 3: Overview of Foundation Model based PIS (§5).
@@ 5 Foundation Model based PIS
As shown in Fig. 3, this section reviews FM-based PIS methods.
@@ 5.1.1 SAM-based Solution
$\bullet$ Multi-Granularity Interactive Segmentation. Most existing interactive segmentation methods determines a single segmentation mask based on users’ input, which ignores spatial ambiguity. In contrast, SAM introduces a multi-granularity interactive segmentation pipeline, i.e., for each user interaction, desired segmentation region may be the concept of objects with different parts nearby. To improve the segmentation quality, HQ-SAM [201] proposes a lightweight high-quality output token replace the original SAM’s output token. After training on 44K highly-accurate masks, HQ-SAM significantly boosts the mask prediction quality of SAM. Since SAM is class-agnostic, a line of works [221, 222] tunes SAM by aligning the query-segmented regions with corresponding textual representations from CLIP. OVSAM [222] explores dual knowledge transfer between SAM and CLIP to enhance SAM’s recognition capabilities. Semantic SAM [217] designs a SAM-like framework that supports multi-granularity segmentation using the captioned SAM data. Although these multi-granularity interactive segmentation approaches alleviate spatial ambiguity, they result in excessive output redundancy and limited scalability. To solve this, GraCo [223] explores granularity-controllable interactive segmentation, which allows precise control of prediction granularity to resolve ambiguity.
@@ 5.2.3 LLMs/MLLMs-based Solution
The success of LLMs/MLLMs has showcased incredible reasoning ability and can answer complex questions, thereby bringing new possibilities to achieve new pixel reasoning and understanding capabilities. In particularly, LISA [59] studies a new segmentation task, called reasoning segmentation. Different from traditional referring segmentation, the segmentors in this setting are developed to segment the object based on implicit query text involving complex reasoning. Notably, the query text is not limited to a straightforward reference (e.g., “the front-runner”), but a a more complicated description involving complex reasoning or world knowledge (e.g., “who will win the race?”). LISA employs LLaVA [251] to output a text response based on the input image, text query, and a [seg] token. The embedding for the customized [seg] token is decoded into the segmentation mask via SAM decoder. Afterwards, LISA++ [252] promotes LISA to differentiate individuals within the same category and enables more natural conversation in multi-turn dialogue. Based on these works, many efforts have been devoted to promote the reasoning capability and segmentation accuracy. LLM-Seg [253] proposes using SAM to generate a group of mask proposals that selects the best-suited answer as the final segmentation prediction. Next-Chat [254] adds a [trigger] token that depicts the coordinate of the object box as a supplementary input for MLLM to help generate better masks. Similarly, GSVA [255] introduces a rejection token [rej] to relieve the empty-target case where the object referred to in the instructions does not exist in the image, leading to the false-positive prediction. Except for the functional token incorporation, [256, 257] propose using diverse textual descriptions, such as object attribute and part, to enhance the object-text connection for accurate reasoning results. Regarding reasoning costing, PixelLLM [60] introduces a lightweight decoder to reduce the computational cost in the reasoning process. Osprey [258] extends MLLMs by incorporating fine-grained mask regions into language instruction, and delivers remarkable pixel-wise visual understanding capabilities.
@@ 5.3.5 LLM/MLLM-based Solution
Figure 4: Illustration of VPInpainting [304], which solves image segmentation as a visual “fill-in-the-blank” problem. (Figure courtesy of [304])
@@ 5.3.6 In-Context Segmentation
The rapid progress of LLMs leads to an emerging ability to learn in-context from just a few examples [38, 45]. Inspired by this, researchers are exploring a related concept in computer vision called in-context segmentation (ICS). From the perspective of segmenting a query image based on a few supports, ICS can be seen as a sub-task of FSS, but it functions directly on pre-trained models without any task-specific finetuning. Most ICL-emerged LLMs are generative models trained through masked language modeling or next token prediction strategies, leading to efforts in ICS that mimic these self-supervised methods. Pioneering work like VPInpainting [304] approaches visual in-context learning as image inpainting. The architecture is illustrated in Fig. 4. It defines visual prompt as a grid-like single image containing an input-output example(s) and a query, then trains an inpainting model (via MAE [305]) to predict the missing parts of an image such that it is consistent with given example(s). With this basis, [306, 307, 308] propose to retrieve optimal examples from large datasets as the prompt. Additionally, Painter [309] and SegGPT [51] are vision generalists built on in-context learning. They unify various vision tasks within the in-context learning framework by standardizing outputs of core vision tasks. Some other studies [310, 311] focus on creating large vision models by formatting images, like language tokenizer, to a group of sequence as visual sentences, and then perform LLM-like training via next token prediction. Notably, developing these visual autoregressive models requires vast amounts of diverse vision data from varied tasks, e.g., image segmentation, depth estimation. PromptDiffusion [312] explores in-context learning for diffusion models by fine-tuning SD to generate the query mask conditioned on the support image-mask pair and the query image. Matcher [313] utilizes DINOv2 to locate the target in query images by bidirectional matching, and leverages the coarse location information as the prompts of SAM for segmentation. Tyche [314] introduces a probabilistic approach to ICS by explicitly modeling training and testing uncertainty, and shows significant potential in medical image segmentation.